Change in Income Change in Consumption Change in Saving
Assumed Increase in Investment $20 $ $4.00
Second Round $ $12.80 $
All other Rounds $ $51.20 $
Totals $ $ $20.00

Refer to the given table, which illustrates the multiplier process. The total change in consumption resulting from the initial change in investment will be:

A) $100

B) $96

C) $180

D) $80
Answer #1

option D)

When Investment increases initially, then total change in Consumption = 20-4 = 16 $

Thus MPC = 16/20 = .8

So in second round , ∆ C = 12.8

In all rounds, ∆ C = 51.2

So total ∆ C = 16+12.8+51.2 = 80